3.9 Show that the commutative ring D is an integral domain if

and only if for a, b, c ∈ D with a #= 0 the relation ab = ac implies that

b = c.


1
Expert's answer
2022-06-22T16:41:40-0400

ab= ac"\implies" ab-ac= 0

"\implies" a(b-c)=0

"\implies" a=0 or b-c=0

"\implies" b-c=0 a≠0 "\implies" b=c